the number N, in millions, of broadband cable and DSL subscribers in the United States t yearsafter 1998 cane be approximated by
N = 0.3t^2 + 0.6t
When will there be 36 million brodband cable and DSL subscribers?
Thank you

Answer by checkley77(12844)

N = 0.3t^2 + 0.6t
36=.3t^2+.6t
.3t^2+.6t-36=0
(.3t-3)(t+12)=0
.3t-3=0
.3t=3
t=3/.3
t=10 years after 1998 or the year 2008 there will be 36 million subscribers.
Proof:
36=.3*10^2+.6^10
36=30+6
36=36